# Towards Characterizing Subjectivity of Individuals through Modeling Value Conflicts and Trade-offs

### v1

@@ Abstract
Large Language Models (LLMs) not only have solved complex reasoning problems but also exhibit remarkable performance in tasks that require subjective decision making. Existing studies suggest that LLM generations can be subjectively grounded to some extent, yet exploring whether LLMs can account for individual-level subjectivity has not been sufficiently studied. In this paper, we characterize subjectivity of individuals on social media and infer their moral judgments using LLMs. We propose a framework, SolAr (Subjective Ground with Value Abstraction), that observes value conflicts and trade-offs in the user-generated texts to better represent subjective ground of individuals. Empirical results show that our framework improves overall inference results as well as performance on controversial situations. Additionally, we qualitatively show that SolAr provides explanations about individuals’ value preferences, which can further account for their judgments.
Towards Characterizing Subjectivity of Individuals through
@@ 1 Introduction
For the last few years, Large Language Models (LLMs) have shifted the paradigm of solving NLP problems to autoregressive language generation and achieved human-like performance in many downstream tasks (Raffel et al., 2020; Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022; Yu et al., 2022; He et al., 2024). Not only have LLMs solved objective problems that require complex reasoning skills, such as STEM-related questions (Imani et al., 2023; Wang et al., 2024c; Abbasiantaeb et al., 2024), they also exhibit remarkable performance in subjective decision-making processes, such as bias detection (Hartvigsen et al., 2022), representing opinions in the survey (Santurkar et al., 2023), generating model evaluation (Perez et al., 2022), social context grounding (Pujari et al., 2024), etc.
@@
Recent studies explore whether LLMs can generate perspectives and reasoning that align well with a specific persona or demographic information (Durmus et al., 2023; Nie et al., 2024; Zheng et al., 2024b). The results of these studies show that it is possible, to an extent, to ground LLM generations with these traits, however, LLMs tend to rely on superficial facts and assumptions about the roles rather than apply a deeper understanding.
@@
Our goal in this paper is to study a different aspect of subjectivity, focusing on the individual-level (rather than generalizing over demographic traits), which has not been sufficiently studied yet. As personalized AI becomes more widely used (McClain, 2024), understanding whether LLMs can be utilized to characterize individual subjectivity becomes more important. Analyzing individual-level subjectivity using LLMs faces two main challenges. The first challenge is guiding LLM generation to be consistent with a specific subjective view. Existing methods for capturing subjectivity at the level of a generalized persona, role, or demographic information show that it is not trivial to steer LLMs to follow certain aspects. Second, even if an optimal approach for grounding subjective aspects in LLMs generations existed, expressing these aspects at the level of an individual is not straightforward (i.e., two instances of the same persona could still have different subjective preferences). Conceptualizing and operationalizing subjectivity at this level poses a second challenge.
@@
Figure 1: Inference process of our framework, SolAr. When an input situation is given, SolAr identifies conflicting values in the story, and retrieve the most relevant value trade-off history from the redditor’s past comments. The retrieved subjective ground is then added to the prompt to infer each redditor’s likely judgments to the situation.
@@
In this paper, our objective is to characterize subjectivity of individuals with LLMs by analyzing users’ behaviors in a Reddit community, r/AmITheAsshole. In this community, original posters write about the situations where they have conflicts with others and ask whether their behaviors are acceptable, and other redditors leave comments with their judgments. Figure 1 shows an example of a post and judgments from different redditors. The research hypothesis is that redditors’ subjective ground, principles that play a fundamental role in human moral judgments (Neuhouser, 1990), can be represented by decomposing their past behaviors; we use redditors’ past comments in the community to determine their likely judgments on unseen situations.
@@
Value pluralism suggests that there are multiple values that may be equally correct, yet in conflict with one another (Crowder, 1998; Galston, 2002). Psychological studies adopt this idea and argue that the human cognitive system makes novel judgments by making trade-offs between conflicting values when it encounters situations with colliding moral intuitions (Fiske and Tetlock, 1997; Guzmán et al., 2022). In this paper, we propose a framework, SolAr (Subjective Ground with Value Abstraction), that observes trade-offs between conflicting values in the user-generated texts, identifies the most relevant value conflicts with respect to the input situation, and infer the redditors’ likely judgments using off-the-shelf LLMs. Our framework aims to tackle the two challenges of characterizing individual-level subjectivity described above; it conceptualizes subjectivity of individuals with trade-offs among abstract values and guides LLM generations with Retrieval-Augmented Generation (RAG). Empirical results suggest that SolAr distinguishes redditors’ subjective preferences and further provides explanations of their value trade-offs. Our proposed framework also improves the overall performance of downstream tasks and better guides LLMs with the prompted subjective ground when it is tested on more controversial situations.
@@
Key Contributions: To the best of our knowledge, this is the first attempt to explore whether off-the-shelf LLMs can be effectively used to account for subjectivity of individuals in the online community. With value abstraction, we highlight that redditors show distinct subjectivity patterns, which makes this research different from social commonsense or opinion mining. Additionally, we propose a novel framework that encompasses a value trade-off system and performs better in downstream tasks as well as grounding LLM generations.
@@ 2 Problem Formulation
Situation Situation refers to the text description of what has happened in the real world. We aim to focus on situations that portray conflicts so that one’s point of view can be projected in diverse ways. “Asking my girlfriend to stay out of apartment for poker night” is an example of situations.
Individual This research aims to analyze different individuals’ reactions and responses on various social situations. Unlike opinion mining or social commonsense studies, where perspectives and rationales are represented as an aggregate of a large number of humans, our main focus is to observe distinct patterns that characterize each individual and understand the rationales behind their decision-making processes.
Subjective Ground Subjective Ground refers to principles or maxims that steer individuals’ moral judgments or perspectives. “One should put their significant other’s needs as top priority.” is a subjective ground item that is relevant to the example situation above. Every individual has a distinct subjective ground and is built from various aspects such as their past experience, demographics, personality, etc. (Eysenck and Eysenck, 1975; Schwaba et al., 2023; Schoeller et al., 2024). One of our research hypotheses is that the subjective ground of individuals can be inferred by observing their past behaviors. Throughout this research, we aim to validate this hypothesis by modeling the subjective ground with individuals’ comment history on social media.
Subjectivity Abstraction One of the limitations of the aforementioned assumption is that it works in an ideal setting in which individuals’ past behaviors can be observed across an extremely large range of social situations—an impractical scenario in real-world contexts. Thus in reality, when inferring individuals’ likely judgments on unseen situations, it is necessary to formulate some hypotheses based on observable subjective ground. Subjectivity Abstraction refers to a process that produces a high-level representation of subjective ground that can be generalized to a broader spectrum of situations. In this paper, we explore several approaches for constructing subjectivity abstraction.
@@ 3 Task Description
We analyze subjective perspectives and judgments of individuals posted on a Reddit community, r/AmITheAsshole. In this community, original posters (OP) describe situations in which they have conflicts with others and ask if they are at fault. Other individuals (i.e., redditors) then leave comments and judge the acceptability of the OP’s behaviors in the situation. We aim to analyze the subjectivity of individuals who leave comments and judgments about the situations.
@@ 3.1 Crawling from r/AmITheAsshole
We crawl all posts in the r/AmITheAsshole community from November 2014 to June 2023 and identify the top 8 redditors who commented the most to ensure a sufficient amount of data for each individual. To ensure that the comments solely mention the situations, we consider top-level comments without including any threaded comments. We then filter out the posts for which the top 8 users did not leave their judgments.
In the r/AmITheAsshole community, redditors leave the judgments in their comments with pre-coded words; YTA (You’re The Asshole), YWBTA (You Would Be The Asshole), NTA (Not The Asshole), YWNBTA (You Would Not Be The Asshole), ESH (Everyone Sucks Here), NAH (No Assholes Here), and INFO (Not Enough Info). For simplicity, we group NTA, NAH, and YWNBTA as ‘acceptable’, and YTA, ESH, and YWBTA as ‘unacceptable’. INFO is discarded as it does not convey subjectivity. Detailed statistics of the crawled dataset are described in Table 1.
Crawled r/AmITheAsshole
# of total instances 18,669
# of unique situations 17,432
Max / Min # of instances per redditor 2,857 / 1,782
# Accept. / Unaccept. labels (overall) 12,279 / 6,390
# Accept. / Unaccept. labels (skewed) 3,251 / 269
Table 1: Statistics of the dataset. The dataset takes the most active 8 redditors into consideration, keeping the instances that contain coded judgments in the comments. Label distributions are described in two ways; by combining all redditors (overall), and by combining two redditors who showed the most skewed judgment patterns (skewed).
3.2 Moral Value Annotation
@@
As discussed in 2, we produce a high-level abstraction of redditors’ comments and the situations. For each pair of situation and an individual’s comment, we prompt off-the-shelf LLMs and generate values that are observed from the situation and redditors’ comments. As human values can be defined and captured differently in the same text, we apply several different approaches.
@@
We first analyze and annotate the texts based on the theory of basic human values proposed by Schwartz (1992). The theory suggests ten basic values that could explain how people in different cultures recognize the underlying motivation and goals. One of the advantages of using this framework is that it explains values that align with or conflict against one another, which naturally expands hypotheses of individual subjectivity. For instance, when we observe an individual’s preference of “Openness to Change”, it is implied that the individual is likely to have less preference of “Conservation” or “Tradition”. Detailed explanations of the ten basic human values and how they are annotated are described in Appendix A.1.
@@
As opposed to using a fixed framework for characterizing human values, we use LLMs to generate more open-ended text of the values observed in the texts. In this setup, we apply value trade-off theories. Humans make judgments based on value trade-offs when different values conflict to each other (Fiske and Tetlock, 1997; Leyva, 2019; Guzmán et al., 2022). As situations in the dataset mostly describe conflicts OPs are having, we prompt LLMs to identify all conflicting value pairs in the situation and discover value trade-offs made by each redditor in the comments.
@@
When conflicting values are generated in an open-ended manner by LLMs, the level of abstraction is insufficient; values describe situation-specific details (e.g. “prioritizing girlfriend’s plan to cook together”), rather than general concepts (e.g. “partner’s emotional needs”). To address this, we iteratively cluster similar value representations and discover high-level definitions for these clustered values. The initial clusters are formed using HDBSCAN (McInnes et al., 2017), and we later use LLMs to create additional clusters for values that remain uncategorized in the initial clustering phase, resulting in 111 clusters in total. Detailed processes for value generation and clustering are described in Appendix A.2.
@@ 3.3 Learning Problem
Our major focus is to use language models as a reasoning machine to understand the subjective ground of individuals and predict their likely behaviors in unseen instances. More specifically, we formulate a binary classification task where the language models are given with situations and redditors’ most relevant subjective ground, either their past comments or value abstractions, and generate a prediction whether the redditors would judge the OP’s behaviors described in input situations acceptable or not.
@@ 4.1 Subjective Ground Retrieval
In order to infer how the target individual would react to the given input, the most relevant subjective ground needs to be retrieved. We design several heuristics to operationalize the retrieval. First, we could simply select the comments that are left on situations similar to the test situations by computing the pairwise distances between the vector representations of the situations.
While retrieving comments from similar situations, we add another heuristics to explicitly retrieve comments that show different judgment patterns and guarantee that LLMs are prompted with different moral judgments. Suppose that LLMs are prompted to predict whether a redditor judges “not paying for my daughter’s wedding” acceptable or not. In observing the past history, it is possible that this redditor commented to the top 5 most similar situations as “not acceptable”. We then try to find other instances where the redditor says differently. When the redditor judges as “acceptable” to “refusing to contribute to my daughter’s wedding after she cancelled the previous one”, LLMs not only understand the redditor’s general judgment tendencies, but also observe special conditions that would change their usual judgment patterns.
@@
Alternatively to computing the distance between situation representations, we could compute the pairwise distances of the abstract values for all situations. The motivation of using value representations is to retrieve comments from more diverse situations; distance among situation representations yields semantically or topically similar situations only, while value similarity can retrieve situations that are similar not in terms of topics, but in terms of high-level values. Consider an input situation about wedding ceremonies for instance, similarity based on situation representations would result in all wedding related situations, while using representation of abstract values could retrieve situations about social appearance, money, family relationship dynamics, etc.
4.2 Judgment Prediction
Given an input and a set of subjective ground provided by the retrieval process, off-the-shelf LLMs predict the likely moral judgment of a redditor. Prediction language models could be designed as a universal language model; this implies that the difference among individuals mostly comes from their subjective ground, and the output reactions derived from the retrieved subjective ground need to be reasonable in general. In this paper, we use the gpt-4o model (Hurst et al., 2024) to predict the acceptability judgments of individuals.

### v2

# SolAr: Towards Characterizing Subjectivity of Individuals through Modeling Value Conflicts and Trade-offs

@@ Abstract
Large Language Models (LLMs) not only have solved complex reasoning problems but also exhibit remarkable performance in tasks that require subjective decision-making. Existing studies suggest that LLM generations can convey subjectivity to some extent, yet exploring whether LLMs can account for individual-level subjectivity has not been sufficiently studied. In this paper, we characterize the subjectivity of individuals on social media and infer their moral judgments using LLMs. We propose a framework, SolAr (Subjective Ground with Value Abstraction), that observes value conflicts and trade-offs in the user-generated texts to better represent subjective ground of individuals. Empirical results demonstrate that our framework enhances overall inference performance, with notable improvements for users with limited data and in controversial situations. Additionally, we qualitatively show that SolAr provides explanations about individuals’ value preferences, which can further account for their judgments.
SolAr: Towards Characterizing Subjectivity of Individuals through
@@ 1 Introduction
For the last few years, Large Language Models (LLMs) have shifted the paradigm of solving NLP problems to autoregressive language generation and achieved human-like performance in many downstream tasks (Raffel et al., 2020; Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022; Yu et al., 2022; He et al., 2024). Not only have LLMs solved objective problems that require complex reasoning skills, such as STEM-related questions (Imani et al., 2023; Wang et al., 2024c; Abbasiantaeb et al., 2024), they also exhibit remarkable performance in subjective decision-making processes, such as detecting toxicity (Hartvigsen et al., 2022), generating model evaluation (Perez et al., 2022), following ethical principles (Bai et al., 2022), etc.
@@
Recent studies explore whether LLMs can generate perspectives and reasoning that align well with a specific persona or demographic information (Durmus et al., 2023; Nie et al., 2024; Zheng et al., 2024b). The results of these studies show that it is possible, to an extent, to ground LLM generations with these traits, however, LLMs tend to rely on superficial facts and assumptions about the roles and demographic traits rather than apply a deeper understanding.
@@
Our goal in this paper is to study a different aspect of subjectivity, focusing on the individual-level (rather than generalizing over demographic traits), which has not been sufficiently studied yet. As personalized AI becomes more widely used (McClain, 2024), understanding whether LLMs can be utilized to characterize individual subjectivity becomes more important. Analyzing individual-level subjectivity using LLMs faces two main challenges. The first challenge is guiding LLM generation to be consistent with a specific subjective view. Existing methods for capturing subjectivity at the level of a generalized persona, role, or demographic information show that it is not trivial to steer LLMs to follow certain aspects (Durmus et al., 2023; Zheng et al., 2024b). Second, even if an optimal approach for grounding subjective aspects in LLMs’ generations existed, expressing these aspects at the level of an individual is not straightforward (i.e., two instances of the same persona could still have different subjective preferences). Conceptualizing and operationalizing subjectivity at this level poses a second challenge.
@@
Figure 1: Inference process of our framework, SolAr. When an input situation is given, SolAr identifies conflicting values in the story, and retrieves the most relevant value trade-off history from the redditor’s past comments. The retrieved subjective ground is then added to the prompt to infer each redditor’s likely judgments to the situation.
@@
In this paper, our objective is to characterize the subjectivity of individuals with LLMs by analyzing users’ behaviors in a Reddit community, r/AmITheAsshole. In this community, original posters write about situations where they have conflicts with others and ask whether their behaviors are acceptable, and other redditors leave comments with their judgments. We formulate a classification task, predicting whether a specific redditor would judge a test situation acceptable or unacceptable, for evaluating different models’ ability to capture subjectivity. In order to perform the classification task more efficiently, we hypothesize that redditors’ subjective ground, principles that play a fundamental role in making moral judgments on others’ situations (Neuhouser, 1990), can be represented by decomposing their past behaviors; we use redditors’ past comments in the community to determine their likely judgments on unseen situations.
@@
More specifically, we adopt value pluralism when considering redditors’ past comments. Value pluralism suggests that there are multiple values that may be equally correct, yet in conflict with one another (Crowder, 1998; Galston, 2002). Psychological studies adopt this idea and argue that the human cognitive system makes novel judgments by making trade-offs between conflicting values when it encounters situations with colliding moral intuitions (Fiske and Tetlock, 1997; Guzmán et al., 2022). In this paper, we propose a framework, SolAr (Subjective Ground with Value Abstraction), that observes trade-offs between conflicting values in the user-generated texts (i.e. comments), identifies the most relevant value conflicts with respect to the input situation, and infer the redditors’ likely judgments using off-the-shelf LLMs. Our framework aims to tackle the two challenges of characterizing individual-level subjectivity described above; it conceptualizes individual-level subjectivity with value trade-offs, and grounds LLM generations with Retrieval-Augmented Generation (RAG). Empirical results show that SolAr distinguishes redditors’ subjective preferences and further provides explanations of their value trade-offs. Our framework improves the overall performance of downstream tasks and better guides LLMs, especially in low-resource user settings (i.e. users with a limited amount of data) and morally controversial scenarios. Figure 1 shows an example of how SolAr determines the judgment of different redditors given a situation.
@@
Key Contributions: To the best of our knowledge, this is the first attempt to explore whether off-the-shelf LLMs can be effectively used to account for individual-level subjectivity in the real-world online community. With value abstraction, we highlight that redditors show distinct subjectivity patterns. Additionally, we propose a novel framework that encompasses a value trade-off system and performs better in downstream tasks as well as grounding LLM generations.
@@ 2 Problem Formulation
Situation Situation refers to the text description of what has happened in the real world. We focus on situations that portray conflicts so that one’s point of view can be projected in diverse ways. “Asking my girlfriend to stay out of the apartment for poker night” is an example of a situation.
Individual This research aims to analyze different individuals’ reactions and responses to various social situations. Unlike opinion mining or social commonsense reasoning studies, where perspectives and rationales are represented as an aggregate of a large number of humans, our main focus is to observe distinct patterns that characterize each individual and understand the rationales behind their decision-making processes.
Subjective Ground Subjective Ground refers to principles or maxims that steer individuals’ moral judgments or perspectives. “One should put their significant other’s needs as a top priority.” is a subjective ground item that is relevant to the example situation above. Every individual has a distinct subjective ground and is built from various aspects such as their past experience, demographics, personality, etc. (Eysenck and Eysenck, 1975; Schwaba et al., 2023; Schoeller et al., 2024). One of our research hypotheses is that the subjective ground of individuals can be inferred by observing their past behaviors. Throughout this research, we aim to validate this hypothesis by modeling the subjective ground with individuals’ comment history on social media.
Value Abstraction One of the limitations of the aforementioned assumption is that it works in a setting in which individuals’ past behaviors can be observed across an extremely large range of social situations—an impractical scenario in real-world contexts. Thus in reality, when inferring individuals’ likely judgments on unseen situations, it is necessary to formulate some hypotheses based on observable past history. Values, which “guide the selection or evaluation of actions, policies, people, and events”, can serve to induce hypotheses about individuals’ moral judgments (Schwartz, 2012). Value abstraction refers to a process that maps past history (i.e. comments) to a high-level representation of subjectivity that can be generalized to a broader spectrum of situations (i.e. values).
@@ 3 Task Description
We analyze subjective perspectives and judgments of individuals posted on a Reddit community, r/AmITheAsshole. In this community, original posters (OP) describe situations in which they have conflicts with others and ask if they are at fault. Other individuals (i.e. redditors) then leave comments and judge the acceptability of the OP’s behaviors in the situation. Redditors’ judgments are pre-coded words in the comment; YTA (You’re The Asshole), YWBTA (You Would Be The Asshole), NTA (Not The Asshole), YWNBTA (You Would Not Be The Asshole), ESH (Everyone Sucks Here), NAH (No Assholes Here), and INFO (Not Enough Info). For simplicity, we group NTA, NAH, and YWNBTA as ‘acceptable’, and YTA, ESH, and YWBTA as ‘unacceptable’. INFO is discarded as it does not convey subjectivity. We aim to learn the subjectivity of individuals by analyzing redditors’ comments and judgment patterns on various situations.
@@
Truncated r/AmITheAsshole
# of total instances 53,280
# of unique situations 17,432
# of unique redditors 100
Max / Min # of instances per redditor 2,870 / 148
# Accept. / Unaccept. labels (overall) 38,365 / 14,915
# Accept. / Unaccept. labels (skewed) 2,615 / 127
Table 1: Statistics of the truncated dataset we use for training and inference. Label distributions are described in two ways; by combining all redditors (overall), and by combining three redditors who showed the most skewed judgment patterns (skewed).
@@ 3.1 Crawling from r/AmITheAsshole
We crawl all posts in the r/AmITheAsshole community from November 2014 to June 2023 and filter out threaded comments to ensure all comments solely argue the acceptability of the situations. In order to perform training and inference more efficiently, we narrow it down to 1.7K unique reddit posts and 100 redditors who commented on these posts. We discuss how we truncate the data and whether the truncated situations and redditors appropriately represent the population distribution in the Appendix A. Detailed statistics of the truncated dataset are described in Table 1, and the data is released for future reproduction  [^1] https://github.com/younggns/solar.
3.2 Abstract Value Annotation
@@
As discussed in 2, we produce a high-level abstraction of redditors’ comments and the situations. For each pair of situations and an individual’s comments, we prompt an off-the-shelf LLM [^2] We use OpenAI’s gpt-4o-0806 model and generate values that are observed from the situation and redditors’ comments. As human values can be defined and captured differently in the same text, we apply several different approaches.
@@
We first analyze and annotate the texts using a top-down approach based on the theory of basic human values proposed by Schwartz (1992). The theory suggests ten basic values that could explain how people in different cultures recognize the underlying motivation and goals. One of the advantages of using this framework is that it explains values that align or conflict with one another, which naturally expands hypotheses of individual subjectivity. Detailed explanations of the ten basic human values and how they are annotated are described in Appendix B.1.
@@
As opposed to using a fixed set of values for characterizing subjectivity, we use a bottom-up approach to discover more open-ended values observed in the texts. In this setup, we apply value trade-off theories. Humans make judgments based on value trade-offs when different values conflict to each other (Fiske and Tetlock, 1997; Leyva, 2019; Guzmán et al., 2022). As situations in the dataset mostly describe conflicts OPs are having, we prompt LLMs to identify all conflicting value pairs in the situation and discover value trade-offs made by each redditor in the comments.
@@
When conflicting values are generated by LLMs, the level of abstraction is insufficient; values describe situation-specific details (e.g. “prioritizing girlfriend’s plan to cook together”), rather than general concepts (e.g. “partner’s emotional needs”). To address this, we iteratively cluster similar value representations and discover high-level definitions for these clustered values. We follow the approach used in Lam et al. (2024) to derive abstract value representations. Detailed processes for prompting LLMs to generate value trade-offs and clustering values are described in Appendix B.2. With more nuanced representations, we argue that the generated values provide richer contextual information than Schwartz’s values and ultimately help better characterize individual subjectivity. As demonstrated in Appendix B.3, the values produced by the LLM integrate multiple Schwartz values, reflecting the multi-dimensional nature of subjective preferences.
@@ 3.3 Learning Problem
Our main focus is to use language models as a reasoning agent to understand individual-level subjectivity and predict their likely behaviors in unseen instances. More specifically, we formulate a binary classification task where the language models are given situations and redditors’ most relevant subjective ground, either their past comments or value preference, and predict whether the redditors would judge the OP’s behaviors acceptable or not.
@@ 4 Model
Let $\mathcal{U}=\{u_{1},u_{2},\dots,u_{n}\}$ denote the set of all distinct redditors, and $\mathcal{S}=\{s_{1},s_{2},\dots,s_{N}\}$ indicate the set of situations (i.e. reddit posts). When an $i$-th redditor commented on a $j$-th situation, we define their comment as $c_{ij}$ and the acceptability judgment as $y_{ij}$ where $y_{ij}\in\{0,1\}$. More specifically, $\mathcal{C}=\{c_{ij}\ |\ u_{i}\ \text{commented on}\ s_{j}\}$ and $\mathcal{Y}=\{y_{ij}\ |\ u_{i}\ \text{commented on}\ s_{j}\}$. After we annotate the moral values of the situations and comments, we get the values of each situation as $s_{j}^{\mathcal{V}}$ and the values of each comment as $c_{ij}^{\mathcal{V}}$. Using a dynamic embedding model $f_{embed}$, all situations, comments, and values are transformed into vectors [^3] We use OpenAI’s text-embedding-3-large model.. We denote these vectors by bold letters such as $\mathbf{s}_{j}=f_{embed}(s_{j})$.
@@ 4.1 Subjective Ground Retrieval
We first define the redditor history data where the retrieval function searches for the most relevant instances to the input. We keep data separate for each redditor and let $\mathcal{D}_{i}$ denote the history data of an $i$-th redditor. $\mathcal{D}_{i}$ contains representations of the situations and comments, namely $\mathcal{D}_{i}=\{(s_{j},s_{j}^{\mathcal{V}},c_{ij},c_{ij}^{\mathcal{V}},y_{ij})\ |\ u_{i}\ \text{commented on}\ s_{j}\}$.
In order to infer how the target individual would react to a test situation, $x$, we design several heuristics to retrieve the most relevant instances. First, we query the retrieval function with the raw situation representations. Each query consists of a user indicator and a vector representation of a test situation, $\mathbf{x}$. Then the retrieval function outputs $k$ nearest instances with respect to the Euclidean distance between the test situation and situations in the history data $\mathcal{D}$. Let R()R(\ ) be the retrieval function, we define subjective ground retrieval with situations by:
$R(u_{i},\mathbf{x})=\underset{\mathbf{s}_{j}\in\mathcal{D}_{i}}{\mathrm{top}\text{-}k}\;\text{dist}(\mathbf{x},\mathbf{s}_{j})^{-1}$ (1)
where $\text{dist}(\mathbf{x},\mathbf{s}_{j})=\|\mathbf{x}-\mathbf{s}_{j}\|_{2}$
@@
As an alternative to computing the distance between situation representations, we could use abstract values. The motivation for using value representations is to retrieve comments from more diverse cases; distance among situations yields topically similar situations only, while using values can retrieve situations that are relevant in terms of high-level values. Consider, for example, an input situation about wedding ceremonies. While retrieving relevant items with situations provides wedding related instances, using abstract value representations could retrieve situations about social appearance, money, family relationship dynamics, etc. Let $R_{val}(\ )$ be the retrieval function that takes a user indicator and a value representation of a test situation, $\mathbf{x}^{\mathcal{V}}$, we define subjective ground retrieval with values by:
$R_{val}(u_{i},\mathbf{x}^{\mathcal{V}})=\underset{\mathbf{s}_{j}^{\mathcal{V}}\in\mathcal{D}_{i}}{\mathrm{top}\text{-}k}\;\text{dist}(\mathbf{x}^{\mathcal{V}},\mathbf{s}_{j}^{\mathcal{V}})^{-1}$ (2)
4.2 Language Model Prompting
After the retrieval functions provide the most relevant past history of a target redditor to the test situation, we prompt off-the-shelf LLMs to predict the likely judgment of the redditor. Essentially, the LLMs are considered as a general reasoner that accounts for individual subjectivity; LLMs are not trained to represent each individual’s subjectivity, but they perform inference based on the given pieces of specific redditor’s past subjective behaviors.
We add $k$ nearest instances as few-shot examples. In each of the few-shot examples, a short description of the retrieved situation, a redditor’s comment, and their judgment are included. In order to observe the usefulness of the annotated values, we also experiment with adding values that are perceived in the comments in few-shot examples. An example prompt is described below:
Prompt Example with Comments Only You will be given examples of a situation, Person X’s comment on the situation, and Person X’s judgment on the situation (i.e. whether it is acceptable or not).
[Situation] Not babysitting my niece
[Comment] When you talk to your brother ..
[Judgment] Acceptable
Now you will be given a new situation. Based on your understanding of Person X’s judgments on different situations, tell me how Person X would judge the new situation.
[Situation] Canceling a family road trip